If an employee deposits Rs. 2,000 at the end of each year into his company’s plan which pays 7% interest compounded quarterly, how much will he have in the account at the end of 5 years?
1
Expert's answer
2019-05-17T12:45:10-0400

Let number of years = n = 5, annual interest = r = 7% = 0.07, number of compounds = m = 4, deposit = R = 2000.

Value Ak of deposit R after k years Ak = R*(1+r/m)k*m.

Future value of the first deposit An = R*(1+r/m)n*m

Future value of the second deposit An-1 = R*(1+r/m)(n-1)*m

.......

Future value of the last deposit A0= R*(1+r/m)0*m

S = A0 + A1 + .. + An = R*(1+r/m)0*m+ R*(1+r/m)1*m+... + R*(1+r/m)n*m =

(this is sum of geometric series with ratio = (1+r/m)m )

=R*( (1+r/m)(n+1)*m-1 )/( (1+r/m)m-1 )

S = 2000( (1+0.07/4)6*4-1 )/( (1+0.07/4)4-1 ) = 14373.7753523



Answer: 14373.7753523